4. In the circuit shown on the right, the rate at which R1



dissipating electrical energy is 20.0W.



(a) Find R1 and R2.



(b) What is the emf of the battery?



(c) Find the current through both R2 and the 10.0-Ω



resistor.



(d) Calculate the total electrical power consumption in all



the resistors and the electrical power delivered by the battery

1
Expert's answer
2022-03-21T12:47:09-0400

Answer


a)resistance

"R=\frac{\rho l }{A}"


b) emf of batter"E=V+ir"


c) current is given

"I=\frac{V}{R}"


d) electrical power

"P=VI"